Constants Periodic Table Part A A typical home uses approximately 1500 kWh of energy per month If the energy came from a nuclear reaction, what mass would have to be converted to energy per year to meet the energy needs of the home? Express your answer using two significant figures g/year Submit Request Answer
Answer #1

Energy required in 1 year = energy in 1 month * number of month in 1 year

= 1500 kwh * 12

= 18000 kwh

= 18000 * 1000 * 3600 J

= 6.48*10^10 J

Now use:

E = m*c^2

6.48*10^10 J = m * (3.0*10^8)^2

m = 7.2*10^-7 Kg

m = 7.2*10^-4 g

Answer: 7.2*10^-4 g/year